1.Solve the differential equation (2𝐷2 + 5𝐷 + 2)𝑦 = 𝑒-1/ 2 x



2. Solve the differential equation (𝐷3 βˆ’ 3𝐷2 + 4𝐷 βˆ’ 2)𝑦 = π‘π‘œπ‘ π—‘.



1
Expert's answer
2022-03-14T12:35:29-0400

1. The auxiliary equation is, "2m^2 + 5m+2=0".


Solving the equation we get, "m = -\frac{1}{2}, -2"


The complementary function is "\text{C.F} = c_1 e^{-\frac{x}{2}} + c_2 e^{-2x}"


"\begin{aligned}
\text{Particular Integral} &= \frac{1}{2D^2+5D+2}e^{-\frac{x}{2}}\\
&= \frac{1}{2(-\frac{1}{2})^2+5(-\frac{1}{2})+2}e^{-\frac{x}{2}}~~(\text{Replace D by }-\frac{1}{2})\\
&=x \frac{1}{4D+5}e^{-\frac{x}{2}}~~~ (\text{Rule failed in the above step})\\
&=\frac{x e^{-\frac{x}{2}}}{3}~~(\text{Replace D by }-\frac{1}{2})\\
\end{aligned}"

The general solution is "y = c_1 e^{-\frac{x}{2}} + c_{2}e^{-2x} + \dfrac{xe^{-\frac{x}{2}}}{3}"


b. The auxiliary equation is, "m^3-3m^2+4m-2=0." Using synthetic division,


"\begin{array}{c|ccc}
1&1 &-3 & 4 & -2\\
& 0 & ~~1 & -2 & ~~~2\\
\hline
&1 & ~-2 & ~~2 & \mid~0\\
\hline
\end{array}"


"\therefore m^3-3m^2+4m-2 = (m-1)(m^2-2m+2) = 0 \implies m = 1, \\ m^2-2m+2 =0.\\"

Solving the quadratic,


"m = \dfrac{2\pm\sqrt{2^2-4\times1\times2}}{2} = \dfrac{2\pm2i}{2} = 1\pm i" .


Hence the complementary function is,

"\text{C.F} = c_1e^{x} + e^{x}(c_2 \cos x + c_3 \sin x)"

"\begin{aligned}
\text{Particular Integral} &= \frac{1}{D^3 - 3D^2+4D-2}\cos x \\
&= \frac{1}{-D + 3 + 4D - 2}\cos x ~~(\text{Replacing $D^2$ by $-1$})\\
&= \frac{1}{1 + 3D}\cos x\\
&= \frac{1-3D}{1 - 9D^{2}}\cos x~~~~~~~~~~~(\text{Multiplying the conjugate})\\
&= \frac{\cos x - 3D(\cos x)}{10}~~~~~(\text{Replacing $D^2$ by $-1$})\\
&=\frac{\cos x + 3\sin x}{10}\\
\end{aligned}"

The general solution is,


"y = e^x(c_1 + c_{2} \cos x + c_3 \sin x) + \dfrac{\cos x + 3\sin x}{10}"